OpenStudy (jada):

two storage sheds are to have the same area. one is square and one is rectangular. The rectangular shed is 4meters wide and 16meters long. How large is one side of the square shed in meters?

jimthompson5910 (jim_thompson5910):

the area of the rectangular shed is 4*16 = 64 square meters

jimthompson5910 (jim_thompson5910):

use the area of a square formula A = s^2 A = area s = side length to find the side length s (in this case, A = 64)